There is an evolutionary psychology hypothesis that being beautiful has a bigger evolutionary advantage for women than for men, and some researchers have concluded that this should mean that beautiful parents will have more daughters than sons.

Gelman and Weakliem (2009) tested this hypothesis by compiling a list of everyone named as one of People Magazine’s “50 most beautiful people” between 1995 and 2000. They then recorded how many sons and how many daughters each of these “beautiful people” had as of 2007.

For each celebrity, let’s calculate the difference between the number of sons they had and the number of daughters they had. Then let’s calculate the mean and standard deviation of these difference scores. (I have rounded slightly, but the results are the same as they found in the experiment.)


We get:

  • N = 256
  • Mean difference = 0.08
  • Standard deviation of difference scores: 1.16

What is the numerator of the standard error of the difference scores  ?

What is the denominator of the standard error of the difference scores ?

Calculate the standard error of the difference scores
Answer #1

Numerator of the standard error of the difference scores = 1.16

Denominator of the standard error of the difference scores = \sqrt{256} = 16

Standard error of the difference scores = 1.16/16 = 0.0725